OpenStudy (anonymous):

Paris has a 15-page manuscript. The equation below shows the relationship between the total cost of editing 15 pages, T, and the cost per page, c, if she gets the 5% discount.

OpenStudy (anonymous):

Where is the equation, and what's your question?

OpenStudy (anonymous):

15c - 0.05(15c) = T is the equation and these are the plugins 4.00$ , 4.95$ , 5.95$ , 12.00$

OpenStudy (anonymous):

Using the equation, what is the best quality of editing that Paris can get done for a maximum of $171?

OpenStudy (anonymous):

It looks like $12 is the answer. Let c=the cost per page c-.05c=171/15 c=12

OpenStudy (anonymous):

thanks